OpenStudy (anonymous):

Point R divides in the ratio 1 : 3. If the x-coordinate of R is -1 and the x-coordinate of P is -3, what is the x-coordinate of Q?

OpenStudy (anonymous):

A.) -1/3 B.) 5 C.) 6 D.) -9

OpenStudy (anonymous):

Fan + Medal

OpenStudy (anonymous):

What does point R divide? QR or QP?

OpenStudy (anonymous):

QP

OpenStudy (anonymous):

@tom982 help

OpenStudy (anonymous):

If R divides QP with a ratio of 1:3, we know PR:RQ = 1:3. We know PR = 2, so QR = 2*3 = 6. The point Q is 6 more than the point R, so is -1+6=5. B.